# Are supernovae driving turbulence in the solar neighborhood?

### 13 Mar 2026 v1

@@ Abstract
Context. Turbulence plays an important role in shaping the interstellar medium, and strongly influences star formation.
@@
Methods. We compare recent H i line-of-sight velocity observations within a volume of radius 70–500 pc centered on the Sun with a suite of $1$ kpc numerical simulations that include two distinct turbulent drivers: (i) supernova (SN) feedback and (ii) imposed large–scale turbulent forcing. For each simulation, we construct synthetic sky maps that closely mimic the observational one, allowing for a consistent comparison between the simulations and the observational data.
@@ 1 Introduction
The interstellar medium (ISM) is the birthplace of stars; to better understand star formation, it is essential to study the mechanisms that occur in the ISM. Turbulence is one of the main regulators of star formation alongside gravity (Federrath and Klessen, 2012). It plays a dual role, either inducing or opposing gravitational collapse (Mac Low and Klessen, 2004).
@@
Many processes contribute to turbulence in the ISM. However, their relative contributions differ and their importance varies across the different scales (Hennebelle and Falgarone, 2012; Klessen and Glover, 2016). Several studies (e.g. Wada et al., 2002; Bournaud et al., 2010; Fensch et al., 2023) have shown that large–scale gravitational instabilities, driven by self–gravity and galactic rotation, can sustain ISM turbulence even in the absence of stellar feedback. Other large-scale processes, like gas accretion into the galaxy, have also been found to be capable of driving turbulence in the ISM (Klessen and Hennebelle, 2010). Among the different feedback processes, supernova (SN) explosions from massive stars are believed to dominate the energy injection into the ISM (Klessen and Glover, 2016). Each event releases $\sim 10^{51}\,\mathrm{erg}$ in a blast wave, which injects both thermal energy and radial momentum into the surrounding gas. The exact contribution from large driving and supernovae remains controversial and may depend on the environment (Iffrig and Hennebelle, 2015; Walch et al., 2015; Ohlin et al., 2019).
@@
Krumholz et al. (2018) argue that in galaxies with low gas surface densities, stellar feedback alone can drive turbulence, whereas in gas-rich, high-redshift disks, turbulence is mainly sustained by energy injection from large-scale gravitational inflows. A similar result was obtained by Brucy et al. (2020, 2023) using 1 kpc ISM-box simulations. They showed that in gas-rich galaxies, SN alone cannot sustain enough turbulence to maintain SFR close to observed levels. Observations from Bacchini et al. (2020) show that in nearby galaxies with low surface density disks, SNe can sustain turbulence without the need for external large-scale drivers. Roman-Oliveira et al. (2024) explored gas-rich galaxies at a redshift $z\approx 4.5$, and also found that the measured velocity dispersions were consistent with the feedback-dominated branch of the Krumholz et al. (2018) model, while including gravitational instabilities overestimated the turbulence.
@@
On smaller scales, simulations of an SN-driven 250-pc-side ISM box by Padoan et al. (2016) showed that supernova explosions can drive and sustain turbulence within molecular clouds. By comparing their SN-driven cloud sample with observations from the Outer Galaxy Survey, they obtained consistent cloud statistics and dynamical properties. Colman et al. (2022) compared 1 kpc ISM-box simulations with Herschel observations of the Large Magellanic Cloud and found a similar result on cloud scales. However, on scales larger than $\sim$60 pc, supernova feedback alone is not sufficient, and an additional large-scale driver is required.
@@
H i 21-cm observations have long been used to study the kinematics and turbulence of the neutral ISM in nearby galaxies (McClure-Griffiths et al., 2023). Large surveys such as THINGS (Walter et al., 2008) have provided high-resolution maps of H i across the disks of nearby galaxies. Using these data, Tamburro et al. (2009) found systematically higher velocity dispersions, $\sigma_{\mathrm{{H\textsc{i}}}}\simeq 10\pm 2\penalty 10000\ \mathrm{km\,s^{-1}}$, in the inner, actively star-forming regions of galaxies, and lower dispersions, $\sigma_{\mathrm{{H\textsc{i}}}}\simeq 6\penalty 10000\ \mathrm{km\,s^{-1}}$, in the outer H i disks. These results suggest that stellar feedback drives the turbulence in star-forming disks, while large-scale galactic processes maintain a baseline level of turbulence at larger radii.
@@
In the nearby ISM, Soler et al. (2025), from here on S25, combined three-dimensional dust models from Edenhofer et al. (2024) with $\mathrm{H\,\textsc{i}}\;21\text{ cm}$ and $\mathrm{CO}(1\!-\!0)$ emission to construct a spatial map of line-of-sight velocity $(v_{\mathrm{los}}$) for atomic and molecular gas within a region spanning $70\text{--}1250\penalty 10000\ \mathrm{pc}$ in distance and $|b|<5^{\circ}$ around the Sun. From these data, they calculated the mean velocity dispersions in $\mathrm{H\,\textsc{i}}$ and $\mathrm{CO}$ $\sigma_{v,\mathrm{H\,\textsc{i}}}=10.8\,\mathrm{km\,s^{-1}}$ and $\sigma_{v,\mathrm{CO}}=6.6\,\mathrm{km\,s^{-1}}$.
@@
In this work, we compare the observed dispersions of H i $v_{\mathrm{los}}$ with numerical simulations in which turbulence is driven either (i) by supernova (SN) feedback, or (ii) by stochastic Fourier-space forcing, meant to represent turbulent energy injected from the large scales. We also include a comparison with the fiducial MHD-4pc run from the TIGRESS suite (Kim and Ostriker, 2017), where turbulence is also driven by SN feedback. Our goal is to determine which mechanism can reproduce the observed level of turbulence in the local ISM.
@@
The paper is organized as follows: Section 2 presents a brief description of the numerical simulations used; Section 3 describes the analysis pipeline and presents the velocity dispersion comparison; Section 4 summarizes our findings and evaluates whether SN feedback or large-scale driving can reproduce the observed level of turbulence in the local ISM.
@@ 2 Numerical setup
Table 1: Simulation parameters. [^1] TURB runs correspond to simulations with imposed large-scale turbulent forcing, SN runs to supernova–driven simulations (Colman et al., 2025), and TIGRESS-MHD4pc is also SN–driven (Kim and Ostriker, 2017). $f_{\rm rms}$ sets the amplitude of the imposed large-scale turbulent forcing and is expressed in units of $1.46\times 10^{-4}\,\mathrm{km\,s^{-1}\,Myr^{-1}}$. $\Sigma$ is the initial gas column density. $B_{0}$ is the initial magnetic-field strength. Coarse resolution and Finest resolution give the coarsest and finest grid spacings. Time indicates the interval over which the synthetic maps are constructed. $\sigma_{v,\mathrm{los}}$ is the mean LOS velocity dispersion computed from the synthetic maps of each simulation after applying a LOS velocity threshold of $\lvert v_{\mathrm{los}}\rvert\leq 30\penalty 10000\ \mathrm{km\,s^{-1}}$.
@@
In this section, we briefly present the numerical simulations used for comparison with the H i data. The simulations fall into two main families that differ in their mechanism for driving turbulence. Table 1 provides a summary of the main simulation parameters.
Section 2.1 introduces the FRIGG simulations, which include both supernova-driven and large-scale–forced runs, while Section 2.2 presents the supernova-driven MHD-4pc simulation from the TIGRESS suite (Kim and Ostriker, 2017).
@@ 2.1 FRIGG Simulations
FRIGG simulations consist of a suite of three-dimensional simulations of a cubic $1\penalty 10000\ \mathrm{kpc}$ region of the interstellar medium. They include two distinct families that differ mainly in their turbulence driving mechanism: supernova-driven runs (SN), taken from Colman et al. (2025), and large-scale turbulent–forcing runs (TURB), introduced in this work.
They follow a setup similar to that of the FRIGG simulation presented by Hennebelle (2018) and are performed using the adaptive mesh refinement (AMR) code RAMSES (Teyssier, 2002), which solves the equations of ideal magnetohydrodynamics (MHD) (Fromang et al., 2006). The base grid consists of $256^{3}$ cells, with periodic boundary conditions in the disk plane ($x$ and $y$) and open boundaries in the vertical ($z$) direction.
Each simulation begins with an initial turbulent velocity field generated with random phases and a Kolmogorov power spectrum. Turbulence is then sustained either by supernova feedback (SN runs) or by a stochastic, large-scale forcing applied in Fourier space (TURB runs). Details of the turbulence driving prescriptions are provided in Appendix A.
@@ 2.2 TIGRESS simulation
Our reference model is the MHD-4pc run of the TIGRESS suite (Kim and Ostriker, 2017). It follows a local shearing box that spans $L_{x}=L_{y}=1024\;\mathrm{pc}$ in the disc plane and $L_{z}=4096\;\mathrm{pc}$ vertically. The grid is uniform with 4 pc cell size; boundary conditions are shearing-periodic in the $x$ direction, periodic in $y$, and vacuum in $z$. They impose an initial $k^{-2}$ velocity field for $t=50\;\mathrm{Myr}$, then decay it to zero by $t=100\;\mathrm{Myr}$, to prevent an artificial collapse burst. After that, turbulence is driven self‐consistently by clustered and runaway supernovae together with time‐dependent FUV heating; no external forcing is applied. The simulation reaches a statistical steady state after $\sim 100\;\mathrm{Myr}$ and is evolved for a total of $700\;\mathrm{Myr}$. More information about this simulation can be found in their paper.
For our analysis, we use the publicly available snapshots [^2] https://princetonuniversity.github.io/astro-tigress/, which consist of a $1024^{3}\;\mathrm{pc^{3}}$ sub‐volume centered on the mid-plane and cover the interval $t\simeq 200\text{–}400\;\mathrm{Myr}$.
@@ 3 Results
Figure 1: Probability density functions of mass-weighted line-of-sight velocity dispersions for all simulations. Each curve shows the distribution of dispersion values obtained from the different masked sky maps for a given simulation. Dashed curves correspond to the large-scale–forced TURB runs, while dotted curves denote the supernova-driven simulations (SN series and TIGRESS-MHD4pc). The dot-dashed vertical line marks the median H i velocity dispersion, $\sigma_{50}=11.1^{+0.9}_{-0.3}\penalty 10000\ \mathrm{km\,s^{-1}}$, derived from the Monte Carlo sampling of the line-of-sight velocities reported by S25, using the Reid et al. (2019) Galactic rotation model parameters in a volume spanning 70–500 pc around the Sun. The gray shaded vertical band represents the corresponding 25th–75th percentile range. Velocity dispersions are computed after applying a velocity threshold of $|v_{\mathrm{los}}|\leq 30\penalty 10000\ \mathrm{km\,s^{-1}}$.
In this section, we compare the results from the different numerical simulations mentioned in Section 2 with the $\mathrm{H\,\textsc{i}}$ line–of–sight velocity statistics reported by S25.
3.1 On the dispersions measured by observation
S25 measured line-of-sight H i velocities in the solar neighborhood by combining three-dimensional dust density models from Edenhofer et al. (2024) with H i line emission. This approach allowed them to reconstruct the local H i velocity field. However, the initial reconstruction includes velocity components that may not be physically associated with the local ISM; instead they may arise from chance correlations in the method or from non-local high-velocity gas along the line of sight. Motivated by the LOS velocities expected in the solar neighborhood from the Galactic rotation model of Reid et al. (2019), they imposed a velocity cut of $|v_{\mathrm{los}}|\leq 25\penalty 10000\ \mathrm{km\,s^{-1}}$ on the input H i spectra to mitigate spurious correlations and exclude non-local emission.
They then subtract the Galactic rotation model from the reconstructed velocities. In the following, when referring to observational line-of-sight velocities, we always subtract the Galactic rotation model. After subtracting it, the resulting LOS velocities can exceed the original $|v_{\mathrm{los}}|\leq 25\penalty 10000\ \mathrm{km\,s^{-1}}$ cut applied to the input H i spectra. We restrict the comparison with the simulations to the distance range $70\text{--}500\penalty 10000\ \mathrm{pc}$. In this distance range, the maximum LOS velocity after subtracting the model is $\sim 30\penalty 10000\ \mathrm{km\,s^{-1}}$. We therefore adopt $|v_{\mathrm{los}}|\leq 30\penalty 10000\ \mathrm{km\,s^{-1}}$ as the velocity threshold for comparison with the simulations. This choice is further supported by the fact that, up to this threshold, the velocity dispersions measured in the inner and outer Galaxy remain comparable, whereas they increasingly diverge at higher velocity thresholds (see Appendix D).
@@ 3.2 Line-of-sight velocity construction
To ensure a consistent resolution across all simulations, we first resample the simulation outputs onto a regular grid with a resolution of 4 pc before constructing the line-of-sight velocity maps.
Following S25, we define the line-of-sight (LOS) velocity seen by an observer as the projection of the local gas velocity $\mathbf{v}=(v_{x},v_{y},v_{z})$ onto the sight-line vector:
$v_{\mathrm{los}}(\mathbf{r})\;=\;\frac{x\,v_{x}+y\,v_{y}+z\,v_{z}}{r}\;,$ (1)
where $x,y,z$ are the coordinates of the cell relative to the observer, and $r=\sqrt{x^{2}+y^{2}+z^{2}}$.
@@
To minimize the biases that can arise from the choice of a particular position, we place the observer successively at a set of uniformly spaced positions in the disc mid-plane. Specifically, the observer coordinates are chosen as $(x_{\rm obs},y_{\rm obs},0)$, with $x_{\rm obs}$ and $y_{\rm obs}$ taking values between 400-400 and $500\penalty 10000\ \mathrm{pc}$, sampled at intervals of $100\penalty 10000\ \mathrm{pc}$. Because the boxes are strictly periodic in $(x,y,0)$, all cell coordinates are re–wrapped relative to each observer.
@@
The shearing box is periodic only in $y$. We therefore shift the observer along the $y$-axis in steps of $50\penalty 10000\ \mathrm{pc}$, spanning the range $450\leq y_{\rm obs}\leq 500\penalty 10000\ \mathrm{pc}$, and along the vertical direction in steps of $25\penalty 10000\ \mathrm{pc}$ within $25\leq z_{\rm obs}\leq+25\penalty 10000\ \mathrm{pc}$, to increase the statistical sample as $x_{\rm obs}$ is kept fixed to avoid complications associated with shear-periodic boundary conditions. No periodic wrapping is applied in $z$; the restriction on latitude $b$ (see below) ensures that all sight lines remain close to the mid-plane and do not cross the vertical vacuum boundaries.
We keep only cells with gas temperatures in the range $100\penalty 10000\ \mathrm{K}<T<10000\penalty 10000\ \mathrm{K}$ and distances $70\leq r\leq 500\penalty 10000\ \mathrm{pc}$. The dependence on the adopted temperature thresholds is examined in Appendix F. The lower distance limit corresponds to the minimum distance covered by the H i data. After applying these spatial cuts, the remaining cells are binned into $4\penalty 10000\ \mathrm{pc}$ radial bins to match the resolution of the numerical simulations. Positions are then projected into Galactic coordinates $(\ell,b)$, and only those with $|b|\leq 5^{\circ}$ are kept. Finally, the sky is divided into $10^{\circ}\!\times\!10^{\circ}$ longitude–latitude tiles, and mass–weighted averages of the line-of-sight velocity and density are computed for each radial bin and tile.
The H i map have a finer radial resolution (logarithmic bins of $0.4\text{-}2.8\penalty 10000\ \mathrm{pc}$) and incomplete sky coverage, unlike the simulations. For a consistent comparison at similar resolution and sampling, we mapped the H i data onto the same $(r,\ell)$ grid used for the simulations, with linear radial bins spanning $70$ to $500\penalty 10000\ \mathrm{pc}$ of $\simeq 4\penalty 10000\ \mathrm{pc}$ and longitude bins of $10^{\circ}$. Within each $(r,\ell)$ cell we compute the mean and standard deviation of the LOS velocity, and record cells with no data as $\mathrm{NaN}$. We then apply the observational $(r,\ell)$ sampling pattern to the simulation maps (produced from the different observer positions at different time steps), so that only the regions corresponding to observed cells are kept. To minimize possible biases caused by the non-uniform longitudinal sampling of the observations, we repeat this procedure after rotating the observational grid by $0^{\circ}$, $90^{\circ}$, $180^{\circ}$, and $270^{\circ}$ in $\ell$, producing four masked realizations for each simulation map.
In Appendix B, we show the line-of-sight velocity map of H i from S25, together with an example of a full simulation map and the four masked versions obtained after applying the different observational samplings.

### 31 Mar 2026 v2

@@ Abstract
Context. Turbulence plays an important role in shaping the interstellar medium, and it strongly influences star formation.
@@
Methods. We compare recent H i line-of-sight velocity observations within a volume of radius 70–500 pc centered on the Sun with a suite of $1$ kpc numerical simulations that include two distinct turbulent drivers: (i) supernova (SN) feedback and (ii) imposed large-scale turbulent forcing. For each simulation, we constructed synthetic sky maps that closely mimic the observational one, allowing for a consistent comparison between the simulations and the observational data.
@@ 1 Introduction
The interstellar medium (ISM) is the birthplace of stars, and to better understand star formation, it is essential to study the mechanisms that occur in the ISM. Turbulence is one of the main regulators of star formation, alongside gravity (Federrath and Klessen, 2012). It plays a dual role, either inducing or opposing gravitational collapse (Mac Low and Klessen, 2004).
@@
Many processes contribute to turbulence in the ISM. However, their relative contributions differ, and their importance varies across the different scales (Hennebelle and Falgarone, 2012; Klessen and Glover, 2016). Several studies (e.g., Wada et al., 2002; Bournaud et al., 2010; Fensch et al., 2023) have shown that large-scale gravitational instabilities, driven by self-gravity and galactic rotation, can sustain ISM turbulence even in the absence of stellar feedback. Other large-scale processes, such as gas accretion into the galaxy, have also been found to be capable of driving turbulence in the ISM (Klessen and Hennebelle, 2010). Among the different feedback processes, supernova (SN) explosions from massive stars are believed to dominate the energy injection into the ISM (Klessen and Glover, 2016). Each event releases $\sim 10^{51}\,\mathrm{erg}$ in a blast wave, which injects both thermal energy and radial momentum into the surrounding gas. The exact contribution from large driving and SNe remains controversial and may depend on the environment (Iffrig and Hennebelle, 2015; Walch et al., 2015; Ohlin et al., 2019).
@@
Krumholz et al. (2018) argue that in galaxies with low gas surface densities, stellar feedback alone can drive turbulence, whereas in gas-rich high-redshift disks, turbulence is mainly sustained by energy injection from large-scale gravitational inflows. A similar result was obtained by Brucy et al. (2020, 2023) using 1 kpc ISM-box simulations. They showed that in gas-rich galaxies, the SN alone cannot sustain enough turbulence to maintain the star formation rate close to observed levels. Observations from Bacchini et al. (2020) showed that in nearby galaxies with low surface density disks, SNe can sustain turbulence without the need for external large-scale drivers. Roman-Oliveira et al. (2024) explored gas-rich galaxies at a redshift of $z\approx 4.5$ and found that the measured velocity dispersions were consistent with the feedback-dominated branch of the Krumholz et al. (2018) model. However, they found that including gravitational instabilities overestimated the turbulence.
@@
On smaller scales, simulations of an SN-driven 250-pc-side ISM box by Padoan et al. (2016) showed that SN explosions can drive and sustain turbulence within molecular clouds. By comparing their SN-driven cloud sample with observations from the Outer Galaxy Survey, they obtained consistent cloud statistics and dynamical properties. Colman et al. (2022) compared 1 kpc ISM-box simulations with Herschel observations of the Large Magellanic Cloud and found a similar result on cloud scales. However, on scales larger than $\sim$60 pc, SN feedback alone is not sufficient, and an additional large-scale driver is required.
@@
We note that H i 21-cm observations have long been used to study the kinematics and turbulence of the neutral ISM in nearby galaxies (McClure-Griffiths et al., 2023). Large surveys such as THINGS (Walter et al., 2008) have provided high-resolution maps of H i across the disks of nearby galaxies. Using these data, Tamburro et al. (2009) found systematically higher velocity dispersions ($\sigma_{\mathrm{{H\textsc{i}}}}\simeq 10\pm 2\penalty 10000\ \mathrm{km\,s^{-1}}$) in the inner actively star-forming regions of galaxies and lower dispersions ($\sigma_{\mathrm{{H\textsc{i}}}}\simeq 6\penalty 10000\ \mathrm{km\,s^{-1}}$) in the outer H i disks. These results suggest that stellar feedback drives the turbulence in star-forming disks, while large-scale galactic processes maintain a baseline level of turbulence at larger radii.
@@
In the nearby ISM, Soler et al. (2025) (from here on S25) combined three-dimensional dust models from Edenhofer et al. (2024) with $\mathrm{H\,\textsc{i}}\;21\text{ cm}$ and $\mathrm{CO}(1\!-\!0)$ emission to construct a spatial map of the line-of-sight (LOS) velocity $(v_{\mathrm{los}}$) for atomic and molecular gas within a region spanning $70\text{--}1250\penalty 10000\ \mathrm{pc}$ in distance and $|b|<5^{\circ}$ around the Sun. From these data, they calculated the mean velocity dispersions in $\mathrm{H\,\textsc{i}}$ and $\mathrm{CO}$: $\sigma_{v,\mathrm{H\,\textsc{i}}}=10.8\,\mathrm{km\,s^{-1}}$ and $\sigma_{v,\mathrm{CO}}=6.6\,\mathrm{km\,s^{-1}}$, respectively.
@@
In this work, we compare the observed dispersions of H i $v_{\mathrm{los}}$ with numerical simulations in which turbulence is driven either by (i) SN feedback or (ii) stochastic Fourier-space forcing, which is meant to represent turbulent energy injected from the large scales. We also include a comparison with the fiducial MHD-4pc run from the TIGRESS suite (Kim and Ostriker, 2017), where turbulence is also driven by SN feedback. Our goal is to determine which mechanism can reproduce the observed level of turbulence in the local ISM.
@@
The paper is organized as follows. Section 2 presents a brief description of the numerical simulations used. In Sect. 3, we describe the analysis pipeline and present the velocity dispersion comparison. In Sect. 4, we summarize our findings and evaluate whether SN feedback or large-scale driving can reproduce the observed level of turbulence in the local ISM.
@@ 2 Numerical setup
Table 1: Simulation parameters. [^1] The TURB runs correspond to simulations with imposed large-scale turbulent forcing, and SN runs correspond to SN-driven simulations (Colman et al., 2025). We note that TIGRESS-MHD4pc is also SN-driven (Kim and Ostriker, 2017). The term $f_{\rm rms}$ sets the amplitude of the imposed large-scale turbulent forcing and is expressed in units of $1.46\times 10^{-4}\,\mathrm{km\,s^{-1}\,Myr^{-1}}$. The term $\Sigma$ is the initial gas column density, and $B_{0}$ is the initial magnetic-field strength. ”Coarse resolution” and ”Finest resolution” give the coarsest and finest grid spacings. ”Time” indicates the interval over which the synthetic maps were constructed. The term $\sigma_{v,\mathrm{los}}$ is the mean LOS velocity dispersion computed from the synthetic maps of each simulation after applying a LOS velocity threshold of $\lvert v_{\mathrm{los}}\rvert\leq 30\penalty 10000\ \mathrm{km\,s^{-1}}$.
@@
In this section we briefly present the numerical simulations used for comparison with the H i data. The simulations fall into two main families that differ in their mechanism for driving turbulence. Table 1 provides a summary of the main simulation parameters. Section 2.1 introduces the FRIGG simulations, which include both SN-driven and large-scale–forced runs, while Sect. 2.2 presents the SN-driven MHD-4pc simulation from the TIGRESS suite (Kim and Ostriker, 2017).
@@ 2.1 FRIGG simulations
The FRIGG simulations consist of a suite of three-dimensional simulations of a cubic $1\penalty 10000\ \mathrm{kpc}$ region of the ISM. The simulations include two distinct families that differ mainly in their turbulence driving mechanism: SN-driven runs (SN), taken from Colman et al. (2025), and large-scale turbulent-forcing runs (TURB), introduced in this work.
These simulations follow a setup similar to that of the FRIGG simulation presented by Hennebelle (2018) and were performed using the adaptive mesh refinement code RAMSES (Teyssier, 2002), which solves the equations of ideal magnetohydrodynamics (MHD; Fromang et al. 2006). The base grid consists of $256^{3}$ cells, with periodic boundary conditions in the disk plane ($x$ and $y$) and open boundaries in the vertical ($z$) direction.
Each simulation begins with an initial turbulent velocity field generated with random phases and a Kolmogorov power spectrum. Turbulence is then sustained either by SN feedback (SN runs) or by stochastic large-scale forcing applied in Fourier space (TURB runs). Details of the turbulence driving prescriptions are provided in Appendix A.
@@ 2.2 TIGRESS simulation
Our reference model is the MHD-4pc run of the TIGRESS suite (Kim and Ostriker, 2017). It follows a local shearing box that spans $L_{x}=L_{y}=1024\;\mathrm{pc}$ in the disk plane and $L_{z}=4096\;\mathrm{pc}$ vertically. The grid is uniform, with a 4 pc cell size. The boundary conditions are shearing-periodic in the $x$ direction, periodic in $y$, and vacuum in $z$. The simulation starts with an initial $k^{-2}$ velocity field for $t=50\;\mathrm{Myr}$ and then decay it to zero by $t=100\;\mathrm{Myr}$ to prevent an artificial collapse burst. After that, turbulence is driven self‐consistently by clustered and runaway SNe together with time‐dependent far-ultraviolet heating; no external forcing is applied. The simulation reaches a statistical steady state after $\sim 100\;\mathrm{Myr}$ and is evolved for a total of $700\;\mathrm{Myr}$. More information about this simulation can be found in their paper. For our analysis, we used the publicly available snapshots, [^2] https://princetonuniversity.github.io/astro-tigress/ which consist of a $1024^{3}\;\mathrm{pc^{3}}$ sub‐volume centered on the mid-plane and cover the interval $t\simeq 200\text{–}400\;\mathrm{Myr}$.
@@ 3 Results
Figure 1: Probability density functions of mass-weighted LOS velocity dispersions for all simulations. Each curve shows the distribution of dispersion values obtained from the different masked sky maps for a given simulation. Dashed curves correspond to the large-scale–forced TURB runs, while dotted curves denote the SN-driven simulations (SN series and TIGRESS-MHD4pc). The dot-dashed vertical line marks the median H i velocity dispersion, $\sigma_{50}=11.1^{+0.9}_{-0.3}\penalty 10000\ \mathrm{km\,s^{-1}}$, derived from the Monte Carlo sampling of the LOS velocities reported by S25 using the Reid et al. (2019) Galactic rotation model parameters in a volume spanning 70–500 pc around the Sun. The gray shaded vertical band represents the corresponding 25th–75th percentile range. Velocity dispersions were computed after applying a velocity threshold of $|v_{\mathrm{los}}|\leq 30\penalty 10000\ \mathrm{km\,s^{-1}}$.
3.1 The dispersions measured by observation
S25 measured LOS H i velocities in the solar neighborhood by combining three-dimensional dust density models from Edenhofer et al. (2024) with H i line emission. This approach allowed them to reconstruct the local H i velocity field. However, the initial reconstruction includes velocity components that may not be physically associated with the local ISM. Instead, they may arise from chance correlations in the method or from nonlocal high-velocity gas along the LOS.
Motivated by the LOS velocities expected in the solar neighborhood from the Galactic rotation model of Reid et al. (2019), S25 imposed a velocity cut of $|v_{\mathrm{los}}|\leq 25\penalty 10000\ \mathrm{km\,s^{-1}}$ on the input H i spectra to mitigate spurious correlations and exclude nonlocal emission. They then subtracted the Galactic rotation model from the reconstructed velocities. In the following, when referring to observational LOS velocities, we always subtract the Galactic rotation model. After subtracting it, the resulting LOS velocities can exceed the original $|v_{\mathrm{los}}|\leq 25\penalty 10000\ \mathrm{km\,s^{-1}}$ cut applied to the input H i spectra.
We restricted the comparison of the simulations to the distance range $70\text{--}500\penalty 10000\ \mathrm{pc}$. In this range, the maximum LOS velocity after subtracting the model is $\sim 30\penalty 10000\ \mathrm{km\,s^{-1}}$. We therefore adopted $|v_{\mathrm{los}}|\leq 30\penalty 10000\ \mathrm{km\,s^{-1}}$ as the velocity threshold for comparison with the simulations. This choice is further supported by the fact that up to this threshold, the velocity dispersions measured in the inner and outer Galaxy remain comparable, whereas they increasingly diverge at higher velocity thresholds (see Appendix D).
@@ 3.2 Line-of-sight velocity construction
To ensure a consistent resolution across all simulations, we first resampled the simulation outputs onto a regular grid with a resolution of 4 pc before constructing the LOS velocity maps. Following S25, we defined the LOS velocity seen by an observer as the projection of the local gas velocity $\mathbf{v}=(v_{x},v_{y},v_{z})$ onto the sight-line vector:
$v_{\mathrm{los}}(\mathbf{r})\;=\;\frac{x\,v_{x}+y\,v_{y}+z\,v_{z}}{r}\;.$ (1)
Here, $x,y,z$ are the coordinates of the cell relative to the observer, and $r=\sqrt{x^{2}+y^{2}+z^{2}}$.
@@
To minimize the biases that can arise from the choice of a particular position, we placed the observer successively at a set of uniformly spaced positions in the disk mid-plane. Specifically, the observer coordinates were chosen as $(x_{\rm obs},y_{\rm obs},0)$, with $x_{\rm obs}$ and $y_{\rm obs}$ taking values between 400-400 and $500\penalty 10000\ \mathrm{pc}$, sampled at intervals of $100\penalty 10000\ \mathrm{pc}$. Because the boxes are strictly periodic in $(x,y,0)$, all cell coordinates are rewrapped relative to each observer.
@@
The shearing box is periodic only in $y$. We therefore shifted the observer along the $y$-axis in steps of $50\penalty 10000\ \mathrm{pc}$, spanning the range $450\leq y_{\rm obs}\leq 500\penalty 10000\ \mathrm{pc}$, and along the vertical direction in steps of $25\penalty 10000\ \mathrm{pc}$ within $25\leq z_{\rm obs}\leq+25\penalty 10000\ \mathrm{pc}$ to increase the statistical sample. In contrast, $x_{\rm obs}$ was kept fixed to avoid complications associated with shear-periodic boundary conditions. No periodic wrapping was applied in $z$. The restriction on latitude $b$ (see below) ensures that all sight lines remain close to the mid-plane and do not cross the vertical vacuum boundaries.
We only kept cells with gas temperatures in the range $100\penalty 10000\ \mathrm{K}<T<10000\penalty 10000\ \mathrm{K}$ and distances $70\leq r\leq 500\penalty 10000\ \mathrm{pc}$. The dependence on the adopted temperature thresholds is examined in Appendix F. The lower distance limit corresponds to the minimum distance covered by the H i data. After applying these spatial cuts, the remaining cells were binned into $4\penalty 10000\ \mathrm{pc}$ radial bins to match the resolution of the numerical simulations. Positions were then transformed into Galactic coordinates $(\ell,b)$, and only those with $|b|\leq 5^{\circ}$ were kept. Finally, the sky was divided into $10^{\circ}\!\times\!10^{\circ}$ longitude-latitude tiles, and mass-weighted averages of the LOS velocity and density were computed for each radial bin and tile.
The H i map has a finer radial resolution (logarithmic bins of $0.4\text{-}2.8\penalty 10000\ \mathrm{pc}$) and incomplete sky coverage, unlike the simulations. For a consistent comparison at a similar resolution and sampling, we mapped the H i data onto the same $(r,\ell)$ grid used for the simulations, with linear radial bins spanning $70$ to $500\penalty 10000\ \mathrm{pc}$ of $\simeq 4\penalty 10000\ \mathrm{pc}$ and longitude bins of $10^{\circ}$. Within each $(r,\ell)$ cell, we computed the mean and standard deviation of the LOS velocity, and we recorded cells with no data as $\mathrm{NaN}$.” We then applied the observational $(r,\ell)$ sampling pattern to the simulation maps (produced from the different observer positions at different time steps) so that only the regions corresponding to observed cells are kept. To minimize any possible biases caused by the nonuniform longitudinal sampling of the observations, we repeated this procedure after rotating the observational grid by $0^{\circ}$, $90^{\circ}$, $180^{\circ}$, and $270^{\circ}$ in $\ell$, thus producing four masked realizations for each simulation map. In Appendix B we show the LOS velocity map of H i from S25 together with an example of a full simulation map and the four masked versions obtained after applying the different observational samplings.